# HuMam: Humanoid Motion Control via End-to-End Deep Reinforcement Learning with Mamba

### v1

@@ Abstract
End-to-end reinforcement learning (RL) for humanoid locomotion is appealing for its compact perception–action mapping, yet practical policies often suffer from training instability, inefficient feature fusion, and high actuation cost. We present HuMam, a state-centric end-to-end RL framework that employs a single-layer Mamba encoder to fuse robot-centric states with oriented footstep targets and a continuous phase clock. The policy outputs joint position targets tracked by a low-level PD loop and is optimized with PPO. A concise six-term reward balances contact quality, swing smoothness, foot placement, posture, and body stability while implicitly promoting energy saving. On the JVRC-1 humanoid in mc-mujoco, HuMam consistently improves learning efficiency, training stability, and overall task performance over a strong feedforward baseline, while reducing power consumption and torque peaks. To our knowledge, this is the first end-to-end humanoid RL controller that adopts Mamba as the fusion backbone, demonstrating tangible gains in efficiency, stability, and control economy.
Index Terms:
Artificial Intelligence, Learning Systems, Cognitive Science, Nonlinear Systems, Cooperative Systems
I Introduction
@@
Humanoid locomotion demands controllers that are both foresightful and resource-aware: foresightful to coordinate accurate foot placement and whole-body balance, and resource-aware to run reliably under onboard compute and actuation limits [1]. End-to-end reinforcement learning (RL) is attractive because it can discover feedback strategies directly from interaction [2]; however, its effectiveness hinges on (i) how heterogeneous inputs are fused and (ii) how training is shaped to avoid trivial or unstable behaviors. Conventional feedforward backbones may under-exploit structure in the observation (e.g., gait phase and oriented footsteps), while heavier sequence models impose unfavorable compute–memory trade-offs [3].
@@
We address these challenges with HuMam, a compact, state-centric RL framework for humanoid walking. HuMam uses a single-layer Mamba encoder as the backbone to fuse (i) robot-centric states—leg joint positions/velocities and base attitude/angular velocity—and (ii) external guidance provided by two oriented planned footsteps (3D position + heading) and a continuous clock that encodes the gait phase. The policy operates at the position level and is executed through a low-gain PD controller, yielding smooth torques while keeping learning well-conditioned [4]. A concise six-term reward promotes contact quality, swing smoothness, precise stepping, upright orientation, nominal height, and upper-body steadiness—encouraging stable, accurate, and energy-aware gaits.
@@
    First Mamba-backed end-to-end humanoid RL controller. We introduced Mamba as a lightweight fusion backbone for humanoid locomotion and show that a single layer suffices to deliver strong performance in a purely state-centric setting.
@@
    Reward shaping for balanced, energy-saving control. A six-term reward jointly regulates foot forces/velocities, stepping accuracy, orientation, height, and upper-body sway, leading to smoother actuation, reduced torque usage, and lower power consumption.
@@
The rest of the paper is structured as follows. Section II surveys reinforcement learning for legged and humanoid locomotion. Section III details our methodology, including the MDP formulation, observation and action spaces, the single-layer Mamba encoder backbone, reward design, and PPO training. Section IV presents the experimental setup and results, with analyses of performance, learning efficiency and stability, energy use, and torque profiles. Section V concludes and outlines directions for future work.
II Related Work
II-A Legged Motion Control Based on Reinforcement Learning
@@
Learning-based locomotion has rapidly progressed from simulation-only proofs of concept to robust, real-world controllers on quadrupeds and humanoids. Pioneering sim-to-real pipelines demonstrated that policies trained purely in simulation with appropriate dynamics randomization can execute agile skills on hardware, enabling precise velocity tracking, fast trotting, and self-righting on ANYmal without hand-crafted gait generators [5]. Subsequent work showed that a blind, purely proprioceptive policy learned via RL can achieve remarkable zero-shot robustness in natural outdoor environments—including rubble, mud, snow, vegetation, and flowing water—by processing streams of joint and base states with a compact neural controller [6].
@@
A complementary line reduces training wall-clock by exploiting massive parallelism in GPU simulators, making on-policy RL practical at scale. By training thousands of agents in parallel with a curriculum over terrains, policies for flat and moderately rough grounds can be obtained in minutes and transferred to the real robot with minimal tuning [7]. Beyond static policies, Rapid Motor Adaptation augments an RL base controller with a learned adaptation module that infers latent environment parameters online, enabling real-time adjustment to novel terrains, payload changes, and wear without explicit system identification [8].
@@
To achieve stable periodic gaits, works such as [9] employed reward shaping based on clock signals to synchronize swing and stance phases, while others [10] leveraged parameterized reference motions from Hybrid Zero Dynamics (HZD) to generate robust walking behaviors. Traversing complex terrain such as stairs has also been demonstrated using purely proprioceptive signals [11], enabling “blind” locomotion that remains robust to unexpected contacts and missteps. While these works highlight the resilience of proprioceptive-only controllers, they also underscore the lack of anticipatory planning, as policies must react after contacts occur. This motivates the integration of exteroceptive modalities, such as planned footsteps, to enable more foresightful locomotion.
II-B Humanoid Motion Control Based on Reinforcement Learning
@@
Applying RL to humanoid platforms poses unique challenges due to larger link masses, high-gear transmissions, and reduced backdrivability, which increase safety risks and make contact-rich exploration more problematic than in quadrupeds or lightweight bipeds. Early progress has therefore been made primarily in high-fidelity simulation. Yang et al. demonstrated Valkyrie gait learning with demonstration-augmented rewards and tracking costs that bias exploration toward humanlike bipedal walking while still allowing RL to discover stabilizing strategies without hand-crafted gait schedules [12]. Similarly, DeepWalk achieved a single omnidirectional walking policy through curriculum scheduling of commanded velocities and contact-aware rewards, producing robust gaits in a simulated humanoid [13].
@@
A second line of work adopts hierarchical designs that combine learned high-level policies with model-based low-level regulators. On the Digit biped, Castillo et al. proposed a cascade structure in which an RL policy outputs task-space setpoints tracked by an operational-space controller, enabling disturbance rejection and transfer to hardware without extensive dynamics randomization [14]. Complementary curriculum-based studies have highlighted the difficulty of precise stepping tasks. The ALLSTEPS framework introduced a curriculum that gradually increases stepping precision and spacing, yielding robust stepping policies for humanoid and Cassie-like robots [15]. More recently, Duan et al. advanced Cassie controllers to closed-loop stepping across challenging stone patterns by mapping footstep commands directly to joint actions, aided by a learned feasibility model and real-time perception [16].
@@
To the best of our knowledge, this is the first work to use Mamba for humanoid locomotion.
III Methodology
III-A Overview
@@
The encoded representation is mapped to policy and value heads that are jointly optimized with proximal policy optimization (PPO). A hierarchical control scheme is adopted, where the high-level policy outputs desired joint positions at 40Hz, while a low-gain proportional–derivative (PD) controller operating at 1000Hz converts them into executable joint torques. The policy naturally incorporates the PD tracking behavior into its predictions, effectively leveraging residual tracking errors as stabilizing interaction forces.
@@
III-B Problem Formulation
1) Markov Decision Process: We model humanoid motion control as a Markov Decision Process (MDP) [2], denoted by $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$. Here, $\mathcal{S}$ represents the state space, $\mathcal{A}$ the action space, $\mathcal{P}(s_{t+1}\mid s_{t},a_{t})$ the transition dynamics governed by the physics simulator, and $r(s_{t},a_{t})$ the immediate reward. The objective is to optimize a policy $\pi_{\theta}(a_{t}\mid s_{t})$ that maximizes the cumulative discounted reward:
@@ III-B Problem Formulation
with coefficients $\alpha_{\text{force}}{=}0.15$, $\alpha_{\text{vel}}{=}0.15$, $\alpha_{\text{step}}{=}0.45$, $\alpha_{\text{orient}}{=}0.05$, $\alpha_{\text{height}}{=}0.05$, and $\alpha_{\text{upper}}{=}0.05$.
@@
III-C Mamba Encoder Backbone
@@
III-D Policy Optimization with PPO
For policy training, we adopt Proximal Policy Optimization (PPO), a widely used on-policy reinforcement learning algorithm well-suited for continuous control tasks. PPO improves stability by constraining policy updates through a clipping mechanism, thereby avoiding abrupt changes that may destabilize training [17]. Formally, the clipped surrogate objective is defined as
@@ III-D Policy Optimization with PPO
IV Experiments
IV-A Environment Setup
All experiments are carried out on a laptop with an 11th Gen Intel(R) Core i9-11900K CPU (16 cores, 3.5 GHz base frequency) and a NVIDIA GeForce RTX 3060 Ti GPU (8 GB, CUDA 12.4). The computing environment is based on Ubuntu 22.04, with physics simulation executed in mc-mujoco [18]. All models are developed in Python 3.12.4 using Ray 2.40.0 and PyTorch 2.4.1. All sampling and training are conducted on the GPU, while multi-thread techniques are adopted on CPU to accelerate the whole learning process in parallel as well.
We conduct experiments on the JVRC-1 humanoid robot, a virtual platform originally designed for the Japan Virtual Robotics Challenge [19]. The model has a height of $1.72$ m and a mass of $62$ kg, and is simulated with full-body dynamics. For locomotion tasks, only the leg joints are actuated, including hip pitch, hip roll, hip yaw, knee, ankle roll, and ankle pitch.
@@ IV-A Environment Setup
We compared our HuMam against Baseline without using the Mamba encoder [20]. Instead, the observations are processed by a conventional feedforward network under identical training settings.
IV-B Implementation Details
@@
As shown in Fig. 1, the policy network consists of a single-layer Mamba encoder for feature fusion, and compact projection heads for policy and value estimation. This design ensures low computational overhead while retaining structured feature interactions. The architectural settings are summarized in Table I.
TABLE I: Network architecture settings.
@@
Projection heads Single-layer MLP, ReLU for policy and value
@@
Curved walking plans are generated using the Humanoid Navigation ROS footstep planner[21], which computes sequences on a 2D occupancy grid given start and goal poses $(x,y,\theta)$. We generate $1000$ trajectories by fixing the start at the origin and sampling goals within $(0,-1,-\pi/2)$ to $(0,1,\pi/2)$, with additional complexity introduced by randomly placed obstacles.
@@
All policies are trained with PPO using on-policy rollouts of length $T$, minibatch updates over multiple epochs, advantage normalization, and gradient clipping. The overall training procedure is outlined in Algorithm 1, and the shared PPO hyperparameters are listed in Table II.
TABLE II: PPO hyperparameters.
@@ IV-C Evaluation Metrics
IV-D Experiment Results
IV-D1 Performance vs. Baseline
Figure 3 compares the learning curves of HuMam and the Baseline across seeds, where solid lines denote the mean return and shaded regions indicate one standard deviation. HuMam consistently achieves higher returns, with a steeper rise at the beginning and a higher plateau in the later stage. Table III further quantifies these differences: HuMam improves peak performance from $269.85$ to $285.50$ (+5.8%) and raises the mean return over the last $10\%$ of training from $263.03$ to $277.50$ (+5.5%). Together, these results demonstrate that HuMam not only achieves a higher performance ceiling but also sustains stronger outcomes in long-horizon training.
@@ IV-D Experiment Results
TABLE III: Comparison of Learning Performance
@@
IV-D2 Effect on Learning Efficiency and Stability
As shown in Figure 3, HuMam exhibits a steeper early slope, higher asymptotic return, and narrower confidence intervals, suggesting faster and more consistent learning.
Quantitative results in Table III further confirm these observations. In terms of efficiency, HuMam requires only 15.815.8M samples to reach a return of $240$, compared to 18.218.2M for the Baseline (13.2% fewer samples). The gap widens at higher thresholds, with HuMam achieving $260$ returns in 20.220.2M samples versus 35.035.0M for the Baseline, a 42.5% improvement.

### v2

@@ Abstract
End-to-end reinforcement learning (RL) for humanoid locomotion is appealing for its compact perception–action mapping, yet practical policies often suffer from training instability, inefficient feature fusion, and high actuation cost. We present HuMam, a state-centric end-to-end RL framework that employs a single-layer Mamba encoder to fuse robot-centric states with oriented footstep targets and a continuous phase clock. The policy outputs joint position targets, tracked by a low-level proportional–derivative controller, and is optimized using proximal policy optimization. A concise six-term reward balances contact quality, swing smoothness, foot placement, posture, and body stability, while implicitly promoting energy efficiency. On the JVRC-1 humanoid in mc-mujoco, HuMam consistently improves learning efficiency, training stability, and overall task performance over a strong feedforward baseline, while reducing power consumption and torque peaks. To our knowledge, this is the first end-to-end humanoid RL controller that adopts Mamba as the fusion backbone, demonstrating tangible gains in efficiency, stability, and control economy. A repository is hosted at https://github.com/allen-legged-robot/humam-rl.
keywords:
[…]
{highlights}
First Mamba-backed humanoid RL controller. Introduces Mamba as a lightweight fusion backbone for end-to-end humanoid locomotion.
Efficient and stable backbone. Mamba enables faster learning, better sample efficiency, and lower variance than baseline.
Energy-aware reward shaping. A six-term PPO reward balances stability, accuracy, smooth actuation, and energy use.
Extensive humanoid validation. Consistent gains on JVRC-1 across diverse locomotion tasks in mc-mujoco.
1 Introduction
Humanoid robots are envisioned as versatile platforms for tasks in human-centric environments, including industrial assistance, healthcare, disaster response, and space exploration [castillo2021digit]. Unlike wheeled or quadrupedal systems, humanoids face heightened challenges: their larger mass, higher center of gravity, and coupled upper–lower body dynamics make locomotion both demanding and safety-critical. Developing reliable locomotion controllers is therefore crucial—not only for advancing embodied intelligence, but also for enabling the safe, efficient, and adaptive deployment of humanoids in real-world scenarios.
@@ 1 Introduction
Humanoid locomotion demands controllers that are both foresightful and resource-aware: foresightful to coordinate accurate foot placement and whole-body balance, and resource-aware to run reliably under onboard compute and actuation limits [tong2024advancements]. End-to-end reinforcement learning (RL) is attractive because it can discover feedback strategies directly from interaction [singh2022reinforcement]; however, its effectiveness hinges on (i) how heterogeneous inputs are fused and (ii) how training is shaped to avoid trivial or unstable behaviors. Conventional feedforward backbones may under-exploit structure in the observation (e.g., gait phase and oriented footsteps), while heavier sequence models impose unfavorable compute–memory trade-offs [saeedvand2019comprehensive].
@@
We address these challenges with HuMam, a compact, state-centric RL framework for humanoid walking. HuMam employs a single-layer Mamba encoder as the backbone to fuse (i) robot-centric states, including leg joint positions and velocities as well as base attitude and angular velocity, and (ii) external guidance specified by two oriented footsteps with three-dimensional position and heading together with a continuous clock that encodes the gait phase. The policy operates at the position level and is executed through a low-gain proportional–derivative (PD) controller, yielding smooth torques while keeping learning well-conditioned [darvish2023teleoperation]. A concise six-term reward promotes contact quality, swing smoothness, precise stepping, upright posture, nominal height, and upper-body stability—encouraging stable, accurate, and energy-efficient gaits.
@@
    First Mamba-backed end-to-end humanoid RL controller. We introduced Mamba as a lightweight fusion backbone for humanoid locomotion and show that a single Mamba layer suffices to deliver strong performance in a purely state-centric setting.
@@
    Reward shaping for balanced, energy-saving control. A six-term reward for proximal policy optimization (PPO) jointly regulates foot contact forces and swing velocities, stepping accuracy, body orientation, torso height, and upper-body sway, leading to smoother actuation, reduced torque usage, and lower power consumption.
@@
The rest of the paper is structured as follows. Section 2 surveys reinforcement learning for legged and humanoid locomotion. Section 3 details our methodology, including the MDP formulation, observation and action spaces, the single-layer Mamba encoder backbone, reward design, and PPO training. Section 4 presents the experimental setup and results, with analyses of performance, learning efficiency and stability, energy use, and torque profiles. Section 5 concludes and outlines directions for future work.
2 Related Work
2.1 Legged Motion Control Based on Reinforcement Learning
@@
Learning-based locomotion has rapidly progressed from simulation-only proofs of concept to robust, real-world controllers on quadrupeds and humanoids. Pioneering sim-to-real pipelines demonstrated that policies trained purely in simulation with appropriate dynamics randomization can execute agile skills on hardware, enabling precise velocity tracking, fast trotting, and self-righting on ANYmal without hand-crafted gait generators [Hwangbo2019Agile]. Subsequent work showed that a blind, purely proprioceptive policy learned via RL can achieve remarkable zero-shot robustness in natural outdoor environments—including rubble, mud, snow, vegetation, and flowing water—by processing streams of joint and base states with a compact neural controller [Lee2020RoughTerrain].
@@
A complementary line reduces training wall-clock by exploiting massive parallelism in GPU simulators, making on-policy RL practical at scale. By training thousands of agents in parallel with a curriculum over terrains, policies for flat and moderately rough grounds can be obtained in minutes and transferred to the real robot with minimal tuning [Rudin2022Minutes]. Beyond static policies, Rapid Motor Adaptation augments an RL base controller with a learned adaptation module that infers latent environment parameters online, enabling real-time adjustment to novel terrains, payload changes, and wear without explicit system identification [Kumar2021RMA].
@@
To achieve stable periodic gaits, works such as [siekmann2021sim] employed reward shaping based on clock signals to synchronize swing and stance phases, while others [li2021reinforcement] leveraged parameterized reference motions from Hybrid Zero Dynamics (HZD) to generate robust walking behaviors. Traversing complex terrain such as stairs has also been demonstrated using purely proprioceptive signals [siekmann2021blind], enabling “blind” locomotion that remains robust to unexpected contacts and missteps. While these works highlight the resilience of proprioceptive-only controllers, they also underscore the lack of anticipatory planning, as policies must react after contacts occur. This motivates the integration of exteroceptive modalities, such as planned footsteps, to enable more foresightful locomotion.
2.2 Humanoid Motion Control Based on Reinforcement Learning
@@
Applying RL to humanoid platforms poses unique challenges due to larger link masses, high-gear transmissions, and reduced backdrivability, which increase safety risks and make contact-rich exploration more problematic than in quadrupeds or lightweight bipeds. Early progress has therefore been made primarily in high-fidelity simulation. [Yang2020WholeBody] demonstrated Valkyrie gait learning with demonstration-augmented rewards and tracking costs that bias exploration toward humanlike bipedal walking while still allowing RL to discover stabilizing strategies without hand-crafted gait schedules. Similarly, DeepWalk achieved a single omnidirectional walking policy through curriculum scheduling of commanded velocities and contact-aware rewards, producing robust gaits in a simulated humanoid [Rodriguez2021DeepWalk].
@@
A second line of work adopts hierarchical designs that combine learned high-level policies with model-based low-level regulators. On the Digit biped, Castillo et al. proposed a cascade structure in which an RL policy outputs task-space setpoints tracked by an operational-space controller, enabling disturbance rejection and transfer to hardware without extensive dynamics randomization [castillo2021robust]. Complementary curriculum-based studies have highlighted the difficulty of precise stepping tasks. The ALLSTEPS framework introduced a curriculum that gradually increases stepping precision and spacing, yielding robust stepping policies for humanoid and Cassie-like robots [Xie2020ALLSTEPS]. More recently, Duan et al. advanced Cassie controllers to closed-loop stepping across challenging stone patterns by mapping footstep commands directly to joint actions, aided by a learned feasibility model and real-time perception [Duan2022SteppingStones].
@@
To the best of our knowledge, this is the first work to introduce Mamba for end-to-end humanoid locomotion with RL.
3 Methodology
3.1 Overview
@@
The encoded representation is mapped to policy and value heads that are jointly optimized with PPO. A hierarchical control scheme is adopted, where the high-level policy outputs desired joint positions at 40Hz, while a low-gain PD controller operating at 1000Hz converts them into executable joint torques. The policy naturally incorporates the PD tracking behavior into its predictions, effectively leveraging residual tracking errors as stabilizing interaction forces.
@@
3.2 Problem Formulation
1) Markov Decision Process: We model humanoid motion control as a Markov Decision Process (MDP) [singh2022reinforcement], denoted by $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$. Here, $\mathcal{S}$ represents the state space, $\mathcal{A}$ the action space, $\mathcal{P}(s_{t+1}\mid s_{t},a_{t})$ the transition dynamics governed by the physics simulator, and $r(s_{t},a_{t})$ the immediate reward. The objective is to optimize a policy $\pi_{\theta}(a_{t}\mid s_{t})$ that maximizes the cumulative discounted reward:
@@ 3.2 Problem Formulation
where $\alpha_{\text{force}}{=}0.15$, $\alpha_{\text{vel}}{=}0.15$, $\alpha_{\text{step}}{=}0.45$, $\alpha_{\text{orient}}{=}0.05$, $\alpha_{\text{height}}{=}0.05$, and $\alpha_{\text{upper}}{=}0.05$.
@@
3.3 Mamba Encoder Backbone
@@
3.4 Policy Optimization with PPO
For policy training, we adopt PPO, a widely used on-policy reinforcement learning algorithm well-suited for continuous control tasks. PPO improves stability by constraining policy updates through a clipping mechanism, thereby avoiding abrupt changes that may destabilize training [schulman2017proximal]. Formally, the clipped surrogate objective is defined as
@@ 3.4 Policy Optimization with PPO
4 Experiments
4.1 Environment Setup
All experiments are carried out on a laptop with an 11th Gen Intel(R) Core i9-11900K CPU (16 cores, 3.5 GHz base frequency) and a NVIDIA GeForce RTX 3060 Ti GPU (8 GB, CUDA 12.4). The computing environment is based on Ubuntu 22.04, with physics simulation executed in mc-mujoco [singh2023mc]. All models are developed in Python 3.12.4 using Ray 2.40.0 and PyTorch 2.4.1. All sampling and training are conducted on the GPU, while multi-thread techniques are adopted on CPU to accelerate the whole learning process in parallel as well.
We conduct experiments on the JVRC-1 humanoid robot, a virtual platform originally designed for the Japan Virtual Robotics Challenge [okugawa2015proposal]. The model has a height of $1.72$ m and a mass of $62$ kg, and is simulated with full-body dynamics. For locomotion tasks, only the leg joints are actuated, including hip pitch, hip roll, hip yaw, knee, ankle roll, and ankle pitch.
@@ 4.1 Environment Setup
We compared our HuMam against Baseline without using the Mamba encoder [singh2022learning]. Instead, the observations are processed by a conventional feedforward network under identical training settings.
4.2 Implementation Details
@@
As shown in Fig. 1, the policy network consists of a single-layer Mamba encoder for feature fusion, and compact projection heads for policy and value estimation. This design ensures low computational overhead while retaining structured feature interactions. The architectural settings are summarized in Table 1.
Table 1: Network architecture settings.
@@
Projection heads Single-layer MLP, ReLU
@@
Curved walking plans are generated using the Humanoid Navigation ROS footstep planner[hornung2012anytime], which computes sequences on a 2D occupancy grid given start and goal poses $(x,y,\theta)$. We generate $1000$ trajectories by fixing the start at the origin and sampling goals within $(0,-1,-\pi/2)$ to $(0,1,\pi/2)$, with additional complexity introduced by randomly placed obstacles.
@@
All policies are trained with PPO using on-policy rollouts of length $T$, minibatch updates over multiple epochs, advantage normalization, and gradient clipping. The overall training procedure is outlined in Algorithm 1, and the shared PPO hyperparameters are listed in Table 2.
Table 2: PPO hyperparameters.
@@ 4.3 Evaluation Metrics
4.4 Experiment Results
4.4.1 Performance vs. Baseline
Fig. 3 compares the learning curves of HuMam and the Baseline across seeds, where solid lines denote the mean return and shaded regions indicate one standard deviation. HuMam consistently achieves higher returns, with a steeper rise at the beginning and a higher plateau in the later stage. Fig. 4, 5, 6, and 7 show the foot trajectories for lateral, backward, forward, and curved walking, respectively. Table 3 further quantifies these differences: HuMam improves peak performance from $269.85$ to $285.50$ (+5.8%) and raises the mean return over the last $10\%$ of training from $263.03$ to $277.50$ (+5.5%). Together, these results demonstrate that HuMam not only achieves a higher performance ceiling but also sustains stronger outcomes in long-horizon training.
@@ 4.4 Experiment Results
Table 3: Comparison of Learning Performance. Mean $\pm$ std over multiple runs.
@@
4.4.2 Effect on Learning Efficiency and Stability
As shown in Fig. 3, HuMam exhibits a steeper early slope, higher asymptotic return, and narrower confidence intervals, suggesting faster and more consistent learning.
Quantitative results in Table 3 further confirm these observations. In terms of efficiency, HuMam requires only 15.815.8M samples to reach a return of $240$, compared to 18.218.2M for the Baseline (13.2% fewer samples). The gap widens at higher thresholds, with HuMam achieving $260$ returns in 20.220.2M samples versus 35.035.0M for the Baseline, a 42.5% improvement.